According to the ideal gas law, if the temperature of a fixed amount of gas rises...

  1. According to the ideal gas law, if the temperature of a fixed amount of gas rises from 300 K (27°C) to 600 K (327°C), what will be its change in volume?  
Answer #1

We know the relationship from gas law equation

V1 / T1 = V2 / T2

V1 / V2 = T1 / T2 ------(1)

V1 is volume of gas at temperature T1

V2 is volume of gas at temperature T2

V1 / V2 = 300 K / 600 K = 0.5

V2 = V1 / 0.5 = 2 V1

Volume of gas will be double.